OpenStudy (anonymous):

PLEASE HELP! The length of a rectangle is 7 1/2 inches and the width is 2 3/4 inches. What is the ratio, using whole numbers, of the length to the width?

OpenStudy (mertsj):

\[\frac{\frac{15}{2}}{\frac{11}{4}}\times\frac{4}{4}=\frac{30}{11}\]